I have to factor some problems out and i can't figure them out.
1. 2x^3-26x^2+80x
2. x^2-15wy+3xy-5wx
3.9x^2-36y^2
4. 3x^2+2x-8
I wouls really appreciate your help.

Answer by Alan3354(69443)

I have to factor some problems out and i can't figure them out.
1. 2x^3-26x^2+80x
= 2x*(x^2 - 13x + 40)
= 2x*(x-5)*(x-8)
------------------
2. x^2-15wy+3xy-5wx
= (x - 5w)*(x + 3y)
--------------
3.9x^2-36y^2
= 9*(x^2 - 4)
= 9*(x-2)*(x+2)
--------------
4. 3x^2+2x-8
= (x + 2)*(3x - 4)